OpenStudy (anonymous):

vector calculus: u=3i+2j-4k and v=i+6j+5k, find scalars a and b so that the vector w=5i-18j-32k can be written in the form w=au+bv

OpenStudy (dumbcow):

set up system of equations 3a +b = 5 2a +6b = -18 -4a +5b = -32 a = 3 b= -4

hero (hero):

Where did you get those equations from?

OpenStudy (anonymous):

<5,-18,-32>=a<3,2,-4>+b<1,6,5>

OpenStudy (anonymous):

3a+b=5

OpenStudy (anonymous):

however i'm not sure if there is infinite solutions in this case but dumbcow has set it up right

OpenStudy (dumbcow):

from what outkast is doing

hero (hero):

And what would that be?

OpenStudy (anonymous):

you could use linear matrix to solve

OpenStudy (anonymous):

basically i'm taking the vector components and equaling them

hero (hero):

Why are you allowed to do that?

OpenStudy (anonymous):

so if you have a<3,2,1> or <v1,v2,v3> , then the scalar of that would be <3a,2a,1a>

hero (hero):

No, first of all, where does a and b even come into play?

OpenStudy (anonymous):

yes cv = <kv1,kv2,kv3>

hero (hero):

We were given u and v with i,j, and k components

OpenStudy (anonymous):

i mean kv = and he's asking in the question what scalars a and b are

hero (hero):

What is kv? kilovolts?

OpenStudy (anonymous):

lol k is scalar

hero (hero):

None of this is making any sense

OpenStudy (anonymous):

if k is a scalar then kv=<kv1,kv2,kv3>

hero (hero):

I don't know what that means

OpenStudy (anonymous):

it means that if you times a vectors by a scalar your components will be times the scalar

OpenStudy (anonymous):

<i,j,k>=<v1,v2,v3>

hero (hero):

Vector Calc 101....In order for me to understand, you'd have to do a complete presentation of it

hero (hero):

But that's okay. I'm going to leave and come back in five minutes. When I come back, I might understand it better than you.

OpenStudy (anonymous):

yeah i'm taking the class right now, and it can be confusing at times.... so many equations

OpenStudy (anonymous):

anyways dumbcows answer is correct, does not have infinite solutions however i wouldn't know how to explain it without using what i've learned in MAT2880

OpenStudy (anonymous):

it tally makes sense, its just multiplying each vectors by its corresponding one. By the way dumcows, wasnet there a shortcut by cross multiplication? or am i confusing that for are of parallelogram

OpenStudy (dumbcow):

shortcut? i dont know maybe, i am not aware of one you could also set up a matrix and solve using reduced row echelon form

OpenStudy (anonymous):

Thank you dumbcow....thats what i did also but i got stuck when i realized i have 3 eqns nd 2 unknowns

hero (hero):

Understanding this from a geometrical standpoint is very helpful